OpenStudy (anonymous):

The driver of a car traveling at 60 ft/sec suddenly applies the brakes. The position of the car is s(t) = 50t − 2t2, t seconds after the driver applies the brakes. How many seconds after the driver applies the brakes does the car come to a stop?

OpenStudy (some.random.cool.kid):

"v = ds/dt or, 60 = d/dt(50t - 2t^2) or, 60 = 50 - 4t or, 4t + 10 = 0 decelaration = f = dv/dt = 4 f/s^2 Time taken to stop = 60/4 = ?" so what should your answer be?

OpenStudy (anonymous):

i'm not sure

OpenStudy (some.random.cool.kid):

Hint: this is the key phrase "Time taken to stop = 60/4 " what would that be?

OpenStudy (anonymous):

oh I only say the equations so it would be 15

OpenStudy (some.random.cool.kid):

yes :D

OpenStudy (some.random.cool.kid):

great job.

OpenStudy (anonymous):

can you help me with another question?